1) You need to administer 125 mg of methylprednisolone (Solu-Medrol) IV push, bid to a patient with acute exacerbation of COPD. You have on hand 40 mg/mL. How much do you prepare?______mL. (round your answer to a whole number)

2) You have an order to give Penicillin 150,000 Units (I.M.) for a patient with strep throat. You have a vial of Penicillin containing 3,000,000 Units in 10ml. How many ml. will you give?_____________
Answer #1

ANSWER.

1) Calculation of ml of drug required.

         Formula = Desired dose/Dose in hand * Quantity.

    Here

        Desired dose (D) = 125 mg.

        Dose in hand (H) = 40 mg.

             Quantity (Q) = 1 ml.

     ml of drug required = D/H * Q

                                    = 125/40 * 1

                                    = 3.12

                                    = 3 ml.

2) Calculation of ml of drug required.

           Formula = Desired dose/Dos in hand * Quantity.

    Here

            Desired dose(D) = 150000 units.

            Dose in hand (H) = 3000000 units.

                Quantity (Q) = 1 ml

As per above formula

            ml of drug required = D/H * Q

                                            = 150000 / 3000000 * 1

                                            = 0.05 * 1

                                            = 0.05 ml.